1) A 1.0 g of protein hemoglobin is dissolved in enough water to make 200 mL of solution. The osmotic pressure of the solution is measured at 25 0C and found to be 1.35 mmHg. What is the molecular weight of the hemoglobin?


2) What is the osmotic pressure of a solution made up of 15 g of salt, NaCl, in 250 mL of solution at 27 0C?


1
Expert's answer
2022-03-04T11:28:04-0500

1) π = iMRT

π - osmotic pressure,

i - van't Hoff's factor,

M - molar concentration of solution (mol/L),

R - ideal gas constant (0.08206 L atm mol-1 K-1),

T - temperature in (K)


π = 1.35 mm Hg = 1.35 mm Hg × (0.001315789 atm / 1 mm Hg) = 0.001776 atm

i = 1

T = 25°C = 25 + 273.15 K = 298.15 K

M = π / (iRT) = 0.001776 atm / (1 × 0.08206 L atm mol-1 K-1 × 298.15 K) = 7.26 × 10-5 mol/L

M = n / V = m / (Mr × V)

Mr = m / (M × V) = 1.0 g / (7.26 × 10-5 mol/L

× 0.200 L) = 68870 g/mol


2) π = iMRT

NaCl = Na+ + Cl-

i = 2

M = n / V = m / (Mr × V) = 15 g / (58.44 g/mol × 0.25 L) = 1.027 mol/L

T = 27°C = 27 + 273.15 K = 300.15 K

π = 2 × 1.027 mol/L × 0.08206 L atm mol-1 K-1 × 300.15 K = 50.59 atm